OpenStudy (anonymous):

A drawer contains 4 red socks, 3 white socks, and 3 blue socks. Without looking, you select a sock at random, replace it, and select a second sock at random. What is the probability that the first sock is blue and the second sock is red?

OpenStudy (kropot72):

The total number of socks = 10 P(blue) = 3/10 P(red) = 4/10 Each selection is independent of the other. Therefore the probability of selecting blue then red is the product of the separate probabilities. Can you now find the probability of the compound event (blue, then red) ?

OpenStudy (anonymous):

I think so, thank you so much! x

OpenStudy (kropot72):

Good! You're welcome :)